# Active Visual Thinking

### v1

@@ Abstract
Complex reasoning problems often involve implicit spatial, geometric, and structural relationships that are not explicitly encoded in text. While recent reasoning models have achieved strong performance across many domains, purely text-based reasoning struggles to represent global structural constraints in complex settings. In this paper, we introduce FigR, which integrates active visual thinking into multi-turn reasoning via end-to-end reinforcement learning. FigR externalizes intermediate structural hypotheses by constructing visual representations during problem solving. By adaptively regulating when and how visual reasoning should be invoked, FigR enables more stable and coherent reasoning over global structural properties that are difficult to capture from text alone. Experiments on challenging mathematical reasoning benchmarks demonstrate that FigR outperforms strong text-only chain-of-thought baselines. In particular, FigR improves the base model by 13.12% on AIME 2025 and 11.00% on BeyondAIME, highlighting the effectiveness of figure-guided multimodal reasoning in enhancing the stability and reliability of complex reasoning.
@@
Active Visual Thinking
@@
Figure 2: Comparison of four reasoning paradigms on a geometry problem with complex spatial constraints. Text-only CoT struggles to capture spatial relations. Unified multimodal CoT often suffers from imprecise visual grounding. Tool-augmented LVLMs are limited by predefined operations. In contrast, FigR generates executable code to construct figures, which enables geometric consistency and leads to correct reasoning.
@@ 1 Introduction
Human reasoning is deeply intertwined with the ability to externalize information into visual form. When faced with complex constraints or multi-step relationships, people naturally draw diagrams, sketch intermediate states, or construct spatial layouts to reduce cognitive load and clarify underlying structure tversky2019diagrams; norman2024things. Cognitive science further shows that such diagrammatic reasoning allows people to discover patterns that are difficult or even impossible to glean from text alone donald1993origins; mandler2010spatial.
@@
Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking illustrates how different reasoning paradigms behave when solving a geometry problem involving angle constraints, tangency conditions, and point intersections. A text-only chain-of-thought (CoT, wei2022chain) model must implicitly track a dense network of spatial relations through symbolic expressions alone. As shown in the leftmost column, this often leads to algebraic errors caused by subtle misinterpretations of positions or angles, since all geometric constraints must be internally maintained without explicit spatial grounding. To alleviate this, a natural extension is to explore unified multimodal models  team2024chameleon; wu2025janus; deng2025emerging; li2025zebra that directly generate images as part of the chain of thought. By externalizing intermediate visual representations, these models aim to make spatial relationships more explicit. However, as illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, such directly generated images are often noisy or spatially imprecise, making it difficult to reliably encode fine-grained geometric constraints. Even minor visual inaccuracies can accumulate during multi-step reasoning and critically impair downstream symbolic computation.
@@
In parallel, a growing line of research investigates tool-augmented large vision-language models (LVLMs), which invoke external visual tools or predefined APIs to assist reasoning, spanning prompt engineering gupta2023visual; hu2024visual; suris2023vipergpt, fine-tuning Wu_2024_CVPR; wang2025pixel, and reinforcement-learning–based approaches zheng2025deepeyes; zhang2025chain. By delegating visual processing to external tools, such methods improve controllability and enforce adherence to predefined operations. Nevertheless, as shown in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, tool-augmented LVLMs remain restricted to a limited set of transformations applied to given images (e.g., zooming or cropping), and cannot autonomously generate task-specific visual representations from scratch. This restriction limits their expressive flexibility when reasoning requires constructing new diagrams to satisfy complex constraints.
@@
Motivated by these limitations, we introduce FigR, which integrates visual construction directly into the reasoning loop through end-to-end reinforcement learning. Instead of generating images directly or relying on fixed toolsets, FigR produces executable code that serves as a bridge between symbolic reasoning and visual rendering. This design enables the model to actively construct and iteratively refine illustrative figures during multi-turn reasoning. The rendered figures then serve as dynamic, interpretable feedback that supports its reasoning process, much like how humans repeatedly sketch and update diagrams when solving complex problems. As illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, this figure-guided process enforces geometric consistency and leads to correct solutions in problems where other paradigms fail. Together, this comparison highlights a key insight: effective visual reasoning requires not only access to visual representations, but also precise, controllable, and interpretable mechanisms for constructing them within the reasoning loop.
@@
To guide this iterative reasoning–rendering process, we design an Adaptive Reward Mechanism that regulates when visual reasoning should be invoked, encouraging selective use of visual cues while discouraging unnecessary or spurious rendering behaviors. This mechanism serves as a control signal that promotes appropriate engagement with figure construction, while accuracy rewards ultimately govern reasoning quality. Crucially, this design eliminates the need for a supervised fine-tuning cold-start stage. Instead, FigR can be initialized directly from an instruction-tuned model (e.g., Qwen3-VL-32B-Instruct bai2025qwen3vl) and refined solely through reinforcement learning. Through this process, FigR autonomously learns when to invoke visual reasoning and how to integrate execution feedback into its reasoning trajectory, without relying on handcrafted heuristics or human-engineered demonstrations.
@@
  • We design a novel adaptive reward mechanism that enables selective use of visual reasoning, without requiring supervised cold-start.
  • We conduct extensive experiments across diverse mathematical reasoning benchmarks, demonstrating substantial gains over text-only baselines.
@@ 2 Related Work
2.1 Text-based Reasoning
@@
A series of works improves upon chain‑of‑thought (CoT) prompting wei2022chain by introducing structured reasoning strategies in large language models (LLMs) achiam2023gpt; touvron2023llama. Program‑of‑Thought (PoT) chen2022program and Chain‑of‑Code li2023chain interleave natural language with executable code so that the models can offload arithmetic and symbolic manipulation to interpreters. Tree‑of‑Thought (ToT) yao2023tree and its variants represent reasoning as a branching search rather than a single chain, allowing models to explore alternative solution paths and prune unpromising branches. DeepSeek-R1 guo2025deepseek successfully applies RL to help models learn when to expand or truncate their reasoning paths. ReTool feng2025retool further teaches models when and how to call a code interpreter during mathematical problem solving, enabling dynamic interleaving of textual reasoning and tool invocation. Nevertheless, even state-of-the-art text-only models struggle with geometry, kinematics, or combinatorial tasks requiring spatial inference. This limitation stems from their inability to externalize intermediate states visually, forcing all constraints to be encoded implicitly in text—exactly the bottleneck illustrated in our introduction.
2.2 Multimodal Reasoning
@@
Large vision–language models (LVLMs) liu2023visual; bai2025qwen2 demonstrated that LLMs can interpret images through relatively shallow alignment layers, but most of them remain limited to text-only outputs. Recent unified architectures team2024chameleon; wu2025janus; deng2025emerging; li2025zebra extend this capability by generating both text and images. However, the quality of the generated images remains limited, preventing them from supporting fine-grained and precise relational depiction. These issues become especially evident in geometric and scientific reasoning tasks, where even minor visual inaccuracies can propagate into significant logical errors.
@@
A parallel literature explores invoking external tools, rendering engines, or predefined APIs as part of the reasoning loop. Recent work demonstrates that LVLMs can decompose a problem, apply visual operators (e.g., zooming, cropping, object detection), and interpret results to refine reasoning gupta2023visual; suris2023vipergpt; hu2024visual; shao2024visual; Wu_2024_CVPR; li2025dyfo; shen2025zoomeye; wang2025pixel. Reinforcement learning–based pipelines zheng2025deepeyes; zhang2025chain further improve tool usage efficiency by rewarding correct invocation patterns. These methods offer strong interpretability and deterministic visual transformations, making them highly effective on perception-centric benchmarks. However, they are inherently restricted to limited operations on given images. They lack support for generating new diagrams, geometric constructions, or abstract visualizations that are not directly present in the input.
@@ 3 Methodology
Figure 3: Overview of our proposed FigR. FigR alternates between textual reasoning and executable drawing steps with visual feedback. An adaptive reward mechanism dynamically regulates visual thinking based on task suitability and execution outcomes.
In this section, we describe the main components of FigR. We first outline the overall architecture and reasoning loop in Section 3.1, then describe the reinforcement-learning (RL) algorithm in Section 3.2, and finally detail the adaptive reward mechanism in Section 3.3. Figure 3 gives an overview of our FigR.
3.1 Figure-Steered Reasoning Loop
Our core idea is to embed visual construction into a model’s reasoning trajectory. Concretely, for each problem input, FigR enters a multi-turn reasoning loop in which it can interleave pure textual reasoning and executable code to generate diagrams - much like a human drawing intermediate sketches while reasoning.
Formally, at each turn $t$, FigR maintains an internal context $h_{t}$ (e.g., textual history, prior code outputs, and rendered images) and optionally a current figure/diagram $I_{t}$. The model’s policy $\pi_{\theta}$ outputs an action:
$a_{t}\sim\pi_{\theta}\left(\cdot\mid h_{t},I_{t}\right).$ (1)
Here, $a_{t}$ may be: (1) A text continuation (e.g., next reasoning step), or (2) An executable code snipper $c_{t}$ written in a general-purpose programming language (e.g., Python). If $a_{t}$ is code, the environment executes it via a sandboxed interpreter, which may produce textual outputs and render a figure:
@@ 3.1 Figure-Steered Reasoning Loop
where $T_{t+1},I_{t+1}$ denotes the textual and visual feedback from code execution, respectively. The context is then updated as:
@@
This multi-turn reasoning repeats until the model emits a special end token (e.g., <End>), or a maximum number of steps is reached. The produced final answer is then evaluated. Overall, the design enables both flexible visual construction and precise code-based control, providing reproducibility and interpretability.
3.2 Reinforcement Learning Procedure
To train the model to decide when to construct figures and how to integrate them into the reasoning process, we employ reinforcement learning. Specifically, we adopt the Group Relative Policy Optimization (GRPO) algorithm shao2024deepseekmath, together with a multi-turn rollout framework implemented in VeRL sheng2024hybridflow.
For each input question, GRPO samples a group of $G$ candidate trajectories (rollouts) under the current policy, computes their final rewards, and uses the group mean reward as a baseline. Trajectories that yield above-average reward are reinforced; below-average ones are suppressed. This yields sample-efficient optimization.
In our adaptation, a state includes: question, previous reasoning and code outputs, all rendered images so far, and any interpreter feedback. Action $a_{t}$ can be a text continuation or a code snippet. The environment handles code execution (if applicable), renders diagrams, and returns observations for the next turn. At the end of each rollout $\tau_{i}$ (i.e., after <End>), we compute the total reward:
$R_{i}=R_{\mathrm{acc}}\left(\tau_{i}\right)+R_{\mathrm{fmt}}\left(\tau_{i}\right)+R_{\mathrm{vis}}\left(\tau_{i}\right),$ (4)
where $R_{\text{acc }}$ measures correctness of the final answer (e.g., 1 if answer matches the ground truth, else 0), $R_{\mathrm{fmt}}$ is the format reward, granted when the model’s output adheres to the required structural format (e.g., a binary signal if the output includes separate reasoning and final answer sections), $R_{\text{vis }}$ regulates when visual reasoning is invoked (details are introduce in Section 3.3).
Within each training batch, for a given prompt, we sample a group of candidate trajectories $\left\{\tau_{i}\right\}_{i=1}^{G}$ using our policy $\pi_{\theta}$. For each trajectory, we compute $R_{i}$. We then compute the group baseline:
$\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i},$ (5)
and the relative advantage for each trajectory:
$\hat{A}_{i}=R_{i}-\bar{R}.$ (6)
GRPO updates the policy by maximizing the following surrogate objective (in expectation):
@@ 3.2 Reinforcement Learning Procedure
Here, $r_{i,t}(\theta)=\frac{\pi_{\theta}\left(a_{i,t}\mid s_{i,t}\right)}{\pi_{\theta_{\text{old }}}\left(a_{i,t}\mid s_{i,t}\right)}$ is the probability ratio; $\varepsilon$ is a clipping hyperparameter; $\pi_{\mathrm{ref}}$ is a reference policy (e.g., initial base policy) used for KL-regularization; $\beta$ is a regularization coefficient.
Above all, this objective encourages the new policy to increase the probability of trajectories whose reward is above the group-average, while penalizing divergence from the reference policy to maintain stability.
@@ 3.3 Adaptive Reward Mechanism
To encourage selective visual thinking while avoiding unnecessary drawing behaviors, we introduce an adaptive reward mechanism. Given an input question $q$, we prompt a larger language model as a classifier to predict a suitability tag $s\in\{0,1\}$, indicating whether drawing-based visual thinking is likely to be beneficial for the question.***In this work, we adopt Deepseek-V3 liu2024deepseek as the classifier because it offers a favorable balance between instruction-following capability and computational cost compared with other tested LLMs. During reinforcement learning, the visual-invocation reward is defined by jointly considering (i) whether the final answer is correct, (ii) whether the problem is suitable for visual reasoning, and (iii) whether the generated code is successfully executed. Let $R_{\mathrm{vis}}$ denote the reward associated with diagram generation, we have:
@@
where $y$ and $\hat{y}$ denote the ground-truth and predicted final answers, respectively, $\mathrm{exec}$ indicates whether the generated code is successfully executed and produces valid visual feedback. Under this formulation, a full visual reward (1.0) is granted only when the model answers correctly and invokes diagram generation appropriately with executable code. A reduced reward (0.2) is assigned when the answer is correct but diagram construction is deemed unnecessary, discouraging redundant visual invocations while still acknowledging successful visual thinking. All other cases, such as incorrect answers or failed executions, will receive zero reward. This design encourages selective and effective use of visual reasoning, while grounding its ultimate utility in final answer correctness.
As a result, the adaptive reward mechanism primarily serves as a control signal that guides when visual reasoning should be invoked. This design reduces unnecessary or spurious figure construction while allowing the model to learn how to integrate diagrammatic feedback effectively into reasoning.

### v2

# Executable Visual States

@@ Abstract
Complex reasoning problems often involve implicit spatial and geometric relationships that are not explicitly encoded in text. While recent reasoning models perform well across many domains, purely text-based reasoning struggles to capture structural constraints in complex settings. In this paper, we introduce FigR, which integrates executable visual construction into multi-turn reasoning via end-to-end reinforcement learning. Rather than relying solely on textual chains of thought, FigR externalizes intermediate hypotheses by generating executable code that constructs diagrams within the reasoning loop. An adaptive reward mechanism selectively regulates when visual construction is invoked, enabling more consistent reasoning over latent global properties that are difficult to infer from text alone. Experiments on eight challenging mathematical benchmarks demonstrate that FigR outperforms strong text-only chain-of-thought baselines, improving the base model by 13.12% on AIME 2025 and 11.00% on BeyondAIME. These results highlight the effectiveness of precise, controllable figure construction of FigR in enhancing complex reasoning ability.
@@
Executable Visual States
@@
Figure 2: Comparison of reasoning paradigms by intermediate state representation. Text-only reasoning relies on implicit symbolic states, unified multimodal models generate uncontrolled visual states, and tool-augmented models operate on given images with predefined transformations. In contrast, FigR constructs executable visual states within the reasoning loop, enabling precise, revisable diagrams that enforce geometric consistency.
@@ 1 Introduction
Human reasoning is deeply intertwined with the ability to externalize information into visual form. When faced with complex constraints or multi-step relationships, people naturally draw diagrams, sketch intermediate states, or construct spatial layouts to reduce cognitive load and clarify underlying structure tversky2019diagrams; norman2024things. Cognitive science further suggests that such diagrammatic reasoning allows people to discover patterns that are difficult or even impossible to glean from text alone donald1993origins; mandler2010spatial.
@@
Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Executable Visual States provides an illustrative comparison of how different reasoning paradigms behave on a geometry problem involving angle constraints, tangency conditions, and point intersections. In text-only chain-of-thought (CoT, wei2022chain) models, all spatial relations must be implicitly maintained through symbolic expressions. This places a heavy burden on internal representations and often leads to cascading algebraic errors when subtle geometric constraints are misinterpreted. A natural extension is to consider unified multimodal models  team2024chameleon; wu2025janus; deng2025emerging; li2025zebra, which generate images as part of the reasoning process to make spatial relationships explicit. While conceptually appealing, these models lack precise control over the generated visual content. Since image generation is not grounded in executable constraints, even small spatial inconsistencies can propagate across reasoning steps, limiting their reliability in problems that require fine-grained geometric precision.
@@
In parallel, tool-augmented large vision-language models (LVLMs) incorporate external visual tools or predefined APIs to assist reasoning, spanning prompt engineering gupta2023visual; hu2024visual; suris2023vipergpt, fine-tuning Wu_2024_CVPR; wang2025pixel, or reinforcement-learning zheng2025deepeyes; zhang2025chain. By delegating visual operations to explicit tools, these approaches improve controllability and execution fidelity. However, they are fundamentally constrained to operating on given images and predefined transformations (e.g., zooming or cropping), which prevents them from autonomously constructing task-specific diagrams required by complex reasoning problems.
@@
Motivated by these limitations, we introduce FigR, which integrates visual construction directly into the reasoning loop via end-to-end reinforcement learning. Rather than relying on unconstrained image generation or fixed toolsets, FigR produces executable code that bridges symbolic reasoning and visual rendering. This allows the model to actively construct and iteratively refine diagrams during multi-turn inference, using rendered figures as explicit, stateful feedback—analogous to how humans repeatedly sketch and update diagrams when solving complex problems. As illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Executable Visual States, this mechanism enforces geometric consistency and supports correct reasoning in cases where other paradigms fail.
@@
To regulate this reasoning–rendering process, we design an Adaptive Reward Mechanism that selectively encourages visual construction when it is beneficial, while discouraging unnecessary or spurious rendering behaviors. Importantly, this design removes the need for a supervised visual reasoning cold-start stage. Instead, FigR can be initialized from an instruction-tuned model (e.g., Qwen3-VL-32B-Instruct bai2025qwen3vl) and refined purely through reinforcement learning. Through this process, FigR autonomously learns both when to invoke visual reasoning and how to integrate execution feedback into its reasoning trajectory, without relying on task-specific supervised visual reasoning data.
@@
  • We introduce an adaptive reward mechanism that regulates the selective use of visual reasoning without requiring supervised cold-start.
  • We demonstrate substantial improvements over strong text-only baselines across diverse mathematical reasoning benchmarks.
@@ 2 Related Work
2.1 Text-based and Programmatic Reasoning
@@
A series of works extends chain-of-thought (CoT) prompting wei2022chain by introducing structured reasoning strategies for large language models (LLMs) achiam2023gpt; touvron2023llama. Program-of-Thought (PoT) chen2022program and Chain-of-Code li2023chain interleave natural language with executable code, allowing models to offload arithmetic and symbolic manipulation to external interpreters. Tree-of-Thought (ToT) yao2023tree and its variants further represent reasoning as a branching search process, enabling exploration of alternative solution paths and pruning of unpromising branches. More recently, reinforcement learning has been applied to regulate reasoning behaviors, such as learning when to expand or truncate reasoning trajectories guo2025deepseek, or when and how to invoke a code interpreter during problem solving feng2025retool.
Despite these advances, such approaches operate primarily over textual or symbolic representations. Intermediate reasoning states remain implicit and must be maintained internally by the model, which poses fundamental challenges for tasks involving geometry, kinematics, or other forms of spatial reasoning. Without explicit external representations, complex relational constraints are prone to the accumulation of errors, limiting the reliability of purely text-based and programmatic reasoning in structurally demanding settings.
2.2 Unconstrained Visual Generation for Reasoning
@@
Large vision-language models (LVLMs) liu2023visual; bai2025qwen2 demonstrate that LLMs can interpret images through relatively shallow alignment layers, but most of them remain limited to text-only outputs. Recent unified multimodal architectures team2024chameleon; wu2025janus; deng2025emerging; li2025zebra further extend generation to both text and images within a single model, enabling visual content to be produced as part of the reasoning process.
While conceptually appealing, these approaches rely on unconstrained image generation without explicit mechanisms to enforce geometric precision or relational consistency. As a result, the generated visual representations are often noisy or spatially imprecise, limiting their effectiveness as reliable intermediate states for reasoning. This limitation becomes particularly pronounced in mathematical and scientific domains, where even minor visual inaccuracies can propagate across reasoning steps and lead to significant downstream errors.
2.3 Tool-Augmented Visual Reasoning
@@
Another line of research augments LVLMs with external visual tools, rendering engines, or predefined APIs. These methods decompose a problem into sub-steps, apply visual operators such as zooming, cropping, or drawing auxiliary lines, and interpret the results to guide subsequent reasoning gupta2023visual; suris2023vipergpt; hu2024visual; shao2024visual; Wu_2024_CVPR; li2025dyfo; shen2025zoomeye; wang2025pixel. Reinforcement learning-based pipelines further improve tool usage efficiency by explicitly rewarding appropriate invocation patterns zheng2025deepeyes; zhang2025chain.
Tool-augmented visual reasoning approaches provide strong controllability, as their visual operations are deterministic and explicitly defined. However, they are inherently restricted to operating on given images and a fixed set of predefined transformations. Consequently, they lack the ability to autonomously construct new diagrams, geometric configurations, or abstract visual representations that are not present in the input. This restriction limits their applicability in reasoning tasks that require dynamic generation and refinement of task-specific visual structures.
@@ 3 Methodology
Figure 3: Overview of FigR. FigR alternates between textual reasoning and executable visual construction within a unified reasoning loop. An adaptive reward mechanism selectively regulates when visual construction is invoked, based on question suitability, execution outcomes, and final answer correctness.
In this section, we describe the main components of FigR. Section 3.1 introduces the executable reasoning loop, Section 3.2 details the reinforcement learning procedure, and Section 3.3 presents the adaptive reward mechanism.
3.1 Executable Visual Construction within the Reasoning Loop
FigR integrates executable visual construction directly into its multi-turn reasoning process. For each input, the model interleaves textual inference and code generation to construct diagrams, analogous to how humans iteratively sketch intermediate states while reasoning.
Formally, at each reasoning step $t$, the model maintains a state consisting of a textual context $h_{t}$ (including prior reasoning steps, code outputs, and interpreter feedback) and an optional rendered diagram $I_{t}$. The policy $\pi_{\theta}$ samples an action:
$a_{t}\sim\pi_{\theta}\left(\cdot\mid h_{t},I_{t}\right),$ (1)
where $a_{t}$ is either (i) a textual continuation or (ii) a code snippet $c_{t}$ generated by the policy itself. When a code action is emitted, it is executed by a sandboxed interpreter, which may produce textual feedback $T_{t+1}$ and/or a rendered diagram $I_{t+1}$:
@@ 3.1 Executable Visual Construction within the Reasoning Loop
The context is then updated accordingly:
@@
This process repeats until a termination token is emitted or a step limit is reached. By representing diagrams as executable and reproducible states, FigR enables precise and controllable visual feedback throughout the reasoning process.
3.2 Reinforcement Learning with Multi-turn Executable Reasoning
Training FigR requires learning long-horizon reasoning policies that interleave textual inference with executable visual construction. In this setting, the quality of individual actions (e.g., whether to emit code at a given step) cannot be evaluated locally, as their utility depends on the final reasoning outcome and the resulting visual states. We therefore formulate learning at the trajectory level and employ reinforcement learning to optimize multi-turn reasoning behaviors.
We adopt Group Relative Policy Optimization (GRPO) shao2024deepseekmath, which is well-suited for trajectory-level supervision with delayed terminal rewards. For each input question, GRPO samples a group of $G$ candidate trajectories under the current policy. Each trajectory $\tau_{i}$ consists of a sequence of textual and code actions, together with their executed interpreter feedback, and terminates when the model emits a special end token. After rollout termination, we compute a scalar reward:
$R_{i}=R_{\mathrm{acc}}(\tau_{i})+R_{\mathrm{fmt}}(\tau_{i})+R_{\mathrm{vis}}(\tau_{i}),$ (4)
which jointly reflects final answer correctness, output format compliance, and the appropriateness of visual construction (Section 3.3). Importantly, rewards are assigned only at the trajectory level, as intermediate actions contribute to reasoning quality through their cumulative effects.
To stabilize learning under such delayed supervision, GRPO compares trajectories generated from the same prompt and uses the group-average reward as a baseline:
$\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}.$ (5)
The relative advantage of each trajectory is then defined as:
$\hat{A}_{i}=R_{i}-\bar{R},$ (6)
which provides a low-variance learning signal that emphasizes relative behavioral differences, such as when executable visual construction leads to better outcomes than purely textual reasoning.
The policy is updated by maximizing the following surrogate objective:
@@ 3.2 Reinforcement Learning with Multi-turn Executable Reasoning
Here, $r_{i,t}(\theta)=\frac{\pi_{\theta}(a_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}(a_{i,t}\mid s_{i,t})}$ denotes the probability ratio between the updated and previous policies, $\varepsilon$ is the clipping threshold, and $\pi_{\mathrm{ref}}$ is a reference policy used for KL regularization. This objective encourages the policy to increase the likelihood of trajectories that achieve higher relative rewards, while maintaining stability during optimization.
Overall, GRPO enables FigR to learn multi-turn reasoning strategies in which the decision to invoke visual construction and the resulting reasoning quality are optimized jointly, without requiring step-wise supervision or an explicit value model.
@@ 3.3 Adaptive Reward Mechanism
While executable visual construction can significantly benefit certain reasoning problems, indiscriminate invocation of diagram generation may introduce unnecessary complexity or distract from effective textual reasoning. To address this trade-off, we introduce an adaptive reward mechanism that acts as a policy-level control signal, guiding when visual construction should be invoked during multi-turn reasoning.
For each input question, we first estimate whether diagrammatic reasoning is likely to be beneficial by querying an auxiliary language model classifier, which outputs a binary suitability label $s\in\{0,1\}$. Importantly, this classifier does not provide step-wise supervision; instead, it modulates reward magnitude, encouraging selective use of visual construction and making the policy robust to moderate misclassification.
During reinforcement learning, the visual-invocation reward $R_{\mathrm{vis}}$ is evaluated only at the trajectory level and is coupled with final answer correctness. Specifically, a positive reward is granted when executable visual construction contributes to a correct solution, with higher weight assigned when diagrammatic reasoning is predicted to be appropriate for the task:
@@
Here, $y$ and $\hat{y}$ denote the ground-truth and predicted final answers, respectively, and $\mathrm{exec}$ indicates successful execution of the generated code.
Rather than explicitly supervising individual drawing actions, this reward formulation biases the policy toward reasoning strategies in which visual construction is used purposefully and effectively. By tying visual rewards to overall task success, the model learns to treat diagram construction as a meaningful intermediate state rather than a mandatory or habitual operation. This design discourages redundant or spurious visual actions while preserving the flexibility to exploit executable visual feedback when it improves reasoning outcomes.